The position of a particle moving along x axis is given by x=6.0t2-1.0t2 where x is in meter and t is in seconds. What is the position of the particle when it achieves its maximum speed is the positive x direction?


1
Expert's answer
2022-01-25T06:49:53-0500

The position of a particle is given by

"x=6.0t^2-1.0t^3"

The speed of particle

"v=(x)'=12t^2-3.0t"

The acceleration

"a=(v)'=24t-3.0"

The particle achieves its maximum speed when "a=0", so

"24t-3.0=0,\quad t=8\:\rm s"

The position of particle at this instant

"x=6.0*8^2-1.0*8^3=-128\:\rm m"